OpenStudy (anonymous):

How would I do this? : 1. Rewrite y = x2 + 14x + 29 in general form

OpenStudy (jackellyn):

That is already in general form

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

unless you mean vertex form?

OpenStudy (anonymous):

this is what my assignment looks like: 1. Rewrite y = x2 + 14x + 29 in general form. (1 point) 2. Rewrite y = 3x2 - 24x + 10 in general form. (1 point) 3. Solve for x: (x - 9)2 = 1 (1 point) 4. Solve for x: x2 + 24x + 90 = 0 (1 point) 5. Solve for x: 2x2 - 4x - 14 = 0 (1 point) so it is not already in that for, it can't be

OpenStudy (jackellyn):

It is, general form is ax2 + bx + c

OpenStudy (jackellyn):

the first two are

OpenStudy (anonymous):

If it already is int he form.. why would it be on there to do?

OpenStudy (jackellyn):

it's a trick question lol

jimthompson5910 (jim_thompson5910):

maybe it wants it in the form 4p(y-k) = (x-h)^2

OpenStudy (anonymous):

Now, how did you get that @jim_thompson5910

OpenStudy (jackellyn):

It says general form.

jimthompson5910 (jim_thompson5910):

Check out this page http://www.purplemath.com/modules/parabola.htm It's towards the bottom

OpenStudy (anonymous):

I'm still lost yall....

OpenStudy (anonymous):

thanks... @jackellyn .

OpenStudy (jackellyn):

NP